# MGSO: Monocular Real-time Photometric SLAM with Efficient 3D Gaussian Splatting

### v2

@@ I Introduction
A key challenge in decoupled systems is the lack of synergy between SLAM and dense reconstruction components. SLAM algorithms often fail to provide optimal data for high-quality dense reconstruction, compromising overall system performance. To address this challenge, we tailored our SLAM system to meet the specific needs of 3D Gaussian Splatting (3DGS) [2]. 3DGS typically requires an initial point cloud to begin reconstruction, with denser, well-structured initial point clouds leading to improved and faster results [3]. However, traditional feature-based SLAM methods produce sparse point clouds that are not optimal for 3DGS initialization. While RGB-D data could provide dense and accurate point clouds, using a monocular camera is preferable for wider applicability.
In this paper, we introduce Monocular-GSO (MGSO), a dense visual SLAM system that performs high-quality online 3D reconstruction in real-time using a single monocular camera. MGSO is a decoupled system that employs photometric SLAM to initialize a 3D Gaussian Splatting (3DGS) module running in parallel, enabling live dense scene reconstruction. The MGSO acronym is a blend of Direct Sparse Odometry (DSO) [4], the photometric SLAM system we built upon, and Gaussian Splatting (GS). In contrast to conventional feature-based SLAM methods that generate sparse point clouds, MGSO is designed to track a dense set of pixels, yielding a denser and well-structured point cloud output. We leverage this dense, structured point cloud to initialize 3D Gaussian Splatting (3DGS) in unmapped areas. Initializing with a high-quality set of points accelerates 3DGS optimization, guiding it toward more compact reconstructions with fewer artifacts and redundancies. As a result, our approach leads to real-time reconstruction with dense 3D maps with high quality and memory compactness.
@@ II Related Work
Dense SLAM research has long explored various 3D representations such as signed distance functions [5], dense point clouds [6], and surfel clouds [7][8]. Despite these advancements, efficiently generating high-quality maps for real-time applications remains challenging. Recent innovations in Neural Radiance Fields (NeRFs) [9] and 3D Gaussian Splatting [2] show promise in addressing the issue. These approaches offer high-quality representations that are easy to create and render. Consequently, this section will focus on systems based on these two techniques.
NeRFs represent scenes using a neural network that outputs novel views based on the input camera’s position and rotation. They also allow for the incremental learning and updating of their 3D representations through gradient-based optimization [9]. This capability has been effectively applied in pioneering works like iMap [10], and further improved by subsequent systems such as NICE-SLAM [11], Orbeez-SLAM [12], and NeRF-SLAM [13]. However, NeRF-based SLAM systems face two notable challenges: they require predefined scene bounds, which is often impractical in exploratory environments; and their implicit scene representations can be difficult to integrate with other systems. These limitations have spurred the adoption of 3DGS as a more suitable dense reconstructor.
@@
Early 3DGS-based SLAM systems, like MonoGS [14], SplaTAM [15], GS-SLAM [16], and Gaussian-SLAM [17], utilize a one-stage approach where tracking and mapping are tightly coupled. This approach introduces a dependency on map refinement before tracking can proceed, which results in slow performance, as shown in Table I. Even newer coupled systems such as CG-SLAM [18], RTG-SLAM [19], and SplatSLAM [20] struggle to run at speeds faster than 20 fps. To allow dense 3DGS-SLAM to operate faster, two-stage systems like Photo-SLAM [21], IG-SLAM [22] and GS-ICP [23] emerged, decoupling the tracking and mapping functions. Furthermore, the majority of current 3DGS systems heavily rely on depth data to perform 3D reconstruction (Table I), making them dependant on RGB-D sensors.
@@
| MonoGS [14] | Coupled | RGB,RGB-D | <5 |
| SplaTAM [15] | Coupled | RGB-D | <5 |
| GS-SLAM [16] | Coupled | RGB-D | >5, <10 |
| Gaussian-SLAM [17] | Coupled | RGB-D | <5 |
| CG-SLAM [18] | Coupled | RGB-D | >15, <20 |
| RTG-SLAM [19] | Coupled | RGB-D | >15, <20 |
| SplatSLAM [20] | Coupled | RGB | <5 |
| GS-ICP [23] | Decoupled | RGB-D | >30 |
| Photo-SLAM [21] | Decoupled | RGB*,RGB-D | >30 |
| IG-SLAM [22] | Decoupled | RGB | >5, <10 |
@@
Our system, MGSO, improves on existing two-stage 3DGS-based SLAM systems while utilizing only RGB data. It operates at 30 fps or higher, a performance matched only by Photo-SLAM and GS-ICP (see Table I). While MGSO is most similar to Photo-SLAM, which combines 3DGS with ORBSLAM3, we address Photo-SLAM’s tendency to create large, memory-inefficient maps. GS-ICP offers exceptional speed but requires depth data to initialize its iterative closest point tracking, whereas our system operates using only RGB data. Unlike IG-SLAM [22], which uses pseudo-depth RGB-D data at the cost of slower performance, MGSO maintains real-time speeds while generating accurate, compact maps.
@@ III-A SLAM module:
The tracking backbone of our system is built upon a lineage of visual SLAM approaches originating from Direct Sparse Odometry (DSO) [4]. DSO’s key innovation is demonstrating that selective pixel sampling for photometric tracking enhances localization accuracy compared to using all pixels in an image. We chose to build upon DSO because we found that its pixel selection strategy also aligns well with initializing 3DGS. DSO tracks a set of pixels across consecutive frames and , optimizing the camera pose () by minimizing the photometric loss equation below for each pixel tracked,
@@
We observed that 3DGS performs better with slightly misplaced initialized points in flat areas than with none at all. Therefore, we implemented an interpolation method that estimates point locations in low-gradient regions based on nearby tracked pixels. Our method employs the Delaunay triangulation algorithm [24] to divide the image into a series of triangles using tracked pixels as vertices. The depth of each interpolated point is calculated as the average depth of the triangle’s vertices, which generally provides accurate results for pixels on flat surfaces. While feature-based systems like Photo-SLAM also interpolate inactive 2D feature points, our method outperforms theirs due to a higher initial point count and by focusing interpolation on flat areas where it’s most accurate, which can be observed in Figure 1.
@@ III-B Dense Reconstruction
MGSO employs 3DGS as its dense reconstruction method. Following the original 3DGS, we map the scene using a set of anisotropic Gaussians . Each Gaussian is modeled with an opacity, rotation, location, scale, and color. We follow Mono-GS’ [14] technique of representing color using RGB instead of spherical harmonics because Mono-GS showed this increased speed for minimal impact on reconstruction quality. We render RGB images of the map using the original differentiable tile-based rasterization introduced in 3DGS.
@@
In order to improve the speed of our system, we employ Gaussian-pyramid based learning introduced in Photo-SLAM [21] to progressively train the Gaussian map. The pyramid helps accelerate training for live video scenarios. A multi-scale Gaussian pyramid is created by repeatedly smoothing and down-sampling ground-truth image captured by the camera. The photometric loss calculation progresses from using the highest pyramid level for in initial iterations to lower levels as training advances. Furthermore, we use an optimized version of the 3DGS CUDA back-end [26] that is faster than the original.
@@
However, we noticed the emergence of floaters when utilizing the 3DGS’s densification and pruning strategies. To mitigate the presence of floaters, we utilized the adaptive control strategies in AbsGS [28]. Thus, as part of our adaptive control strategy we periodically densify Gaussians with high homo-directional view-space position gradients by splitting or cloning them. Large, high-variance Gaussians are split, while small Gaussians in under-reconstructed regions are cloned. Furthermore, we periodically prune Gaussians with low opacity to remove transparent floaters. We use the same splitting and cloning parameters as original 3DGS.
@@ IV Experiments and Discussions
We evaluate MGSO against the latest state-of-the-art 3DGS dense SLAM systems: MonoGS [14], GlORIE-SLAM [20], Splat-SLAM [15], IG-SLAM [22], and Photo-SLAM [21], to demonstrate our system’s combination of high-quality reconstruction, efficient runtime, and compact maps.
@@ IV-A Implementation and Setup
Datasets: Evaluations are done on the sythetic Replica [29] dataset and real-life EuRoC MAV [30] and TUM-RGBD [1] datasets. These datasets are commonly used to evaluate dense SLAM systems.
@@
Experimental Setup: We utilize the default optimization configuration of 3DGS with the exception of adjusting the densification interval to 1000. We configure the SLAM module (DSO [4]) to the default tracking settings. The parameters for increasing SLAM output point density through the inclusion of untracked points were determined through iterative testing.
@@
Evaluation: To ensure a fair comparison, we evaluated our system by inputting the output 3DGS maps and pose estimation data into the original 3DGS rendering and metric scripts. We evaluate our reconstructions with the standard image quality metrics: PSNR, SSIM [25], and LPIPS [31]. Using an third-party evaluation system rather than built-in metrics offers a more realistic assessment, accounting for real-world factors like potential misalignment between the poses and map. Consistent with other dense SLAM systems, we evaluate on every fifth frame. Photo-SLAM’s evaluations are updated using this methodology to ensure consistency. We did ten runs for the Replica and EuRoC dataset and five runs for the TUM-RGBD dataset. Results for other systems were obtained from their respective publications, with the exception of Mono-GS, whose results were sourced from Splat-SLAM [20].

### v3

@@ I Introduction
A key challenge in decoupled systems is the lack of synergy between SLAM and dense reconstruction components. SLAM algorithms often fail to provide optimal data for high-quality dense reconstruction, compromising overall system performance. To address this challenge, we tailored our SLAM system to meet the specific needs of 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians]. 3DGS typically requires an initial point cloud to begin reconstruction, with denser, well-structured initial point clouds leading to improved and faster results [hu2024realtimegaussiansplattingaccelerating]. However, traditional feature-based SLAM methods produce sparse point clouds that are not optimal for 3DGS initialization. While RGB-D data could provide dense and accurate point clouds, using a monocular camera is preferable for wider applicability.
In this paper, we introduce Monocular-GSO (MGSO), a dense visual SLAM system that performs high-quality online 3D reconstruction in real-time using a single monocular camera. MGSO is a decoupled system that employs photometric SLAM to initialize a 3D Gaussian Splatting (3DGS) module running in parallel, enabling live dense scene reconstruction. The MGSO acronym is a blend of Direct Sparse Odometry (DSO) [dso:engel18], the photometric SLAM system we built upon, and Gaussian Splatting (GS). In contrast to conventional feature-based SLAM methods that generate sparse point clouds, MGSO is designed to track a dense set of pixels, yielding a denser and well-structured point cloud output. We leverage this dense, structured point cloud to initialize 3D Gaussian Splatting (3DGS) in unmapped areas. Initializing with a high-quality set of points accelerates 3DGS optimization, guiding it toward more compact reconstructions with fewer artifacts and redundancies. As a result, our approach leads to real-time reconstruction with dense 3D maps with high quality and memory compactness.
@@ II Related Work
Dense SLAM research has long explored various 3D representations such as signed distance functions [izadi2011kinectfusion], dense point clouds [densergd:kerl2013], and surfel clouds [badslam:schops2019][whelan2016elasticfusion]. Despite these advancements, efficiently generating high-quality maps for real-time applications remains challenging. Recent innovations in Neural Radiance Fields (NeRFs) [mildenhall2020nerf] and 3D Gaussian Splatting [kerbl3Dgaussians] show promise in addressing the issue. These approaches offer high-quality representations that are easy to create and render. Consequently, this section will focus on systems based on these two techniques.
NeRFs represent scenes using a neural network that outputs novel views based on the input camera’s position and rotation. They also allow for the incremental learning and updating of their 3D representations through gradient-based optimization [mildenhall2020nerf]. This capability has been effectively applied in pioneering works like iMap [imap:Sucaretal:ICCV2021], and further improved by subsequent systems such as NICE-SLAM [Zhu2022CVPR], Orbeez-SLAM [10160950], and NeRF-SLAM [10341922]. However, NeRF-based SLAM systems face two notable challenges: they require predefined scene bounds, which is often impractical in exploratory environments; and their implicit scene representations can be difficult to integrate with other systems. These limitations have spurred the adoption of 3DGS as a more suitable dense reconstructor.
@@
Early 3DGS-based SLAM systems, like MonoGS [Matsuki:Murai:etal:CVPR2024], SplaTAM [keetha2024splatam], GS-SLAM [yan2023gs], and Gaussian-SLAM [yugay2023gaussianslam], utilize a one-stage approach where tracking and mapping are tightly coupled. This approach introduces a dependency on map refinement before tracking can proceed, which results in slow performance, as shown in Table I. Even newer coupled systems such as CG-SLAM [hu2024cg], RTG-SLAM [peng2024rtg], and SplatSLAM [sandström2024splatslamgloballyoptimizedrgbonly] struggle to run at speeds faster than 20 fps. To allow dense 3DGS-SLAM to operate faster, two-stage systems like Photo-SLAM [hhuang2024photoslam], IG-SLAM [sarikamis2024igslam] and GS-ICP [Ha2024RgbdGS] emerged, decoupling the tracking and mapping functions. Furthermore, the majority of current 3DGS systems heavily rely on depth data to perform 3D reconstruction (Table I), making them dependant on RGB-D sensors.
@@
| MonoGS [Matsuki:Murai:etal:CVPR2024] | Coupled | RGB,RGB-D | <5 |
| SplaTAM [keetha2024splatam] | Coupled | RGB-D | <5 |
| GS-SLAM [yan2023gs] | Coupled | RGB-D | >5, <10 |
| Gaussian-SLAM [yugay2023gaussianslam] | Coupled | RGB-D | <5 |
| CG-SLAM [hu2024cg] | Coupled | RGB-D | >15, <20 |
| RTG-SLAM [peng2024rtg] | Coupled | RGB-D | >15, <20 |
| SplatSLAM [sandström2024splatslamgloballyoptimizedrgbonly] | Coupled | RGB | <5 |
| GS-ICP [Ha2024RgbdGS] | Decoupled | RGB-D | >30 |
| Photo-SLAM [hhuang2024photoslam] | Decoupled | RGB*,RGB-D | >30 |
| IG-SLAM [sarikamis2024igslam] | Decoupled | RGB | >5, <10 |
@@
Our system, MGSO, improves on existing two-stage 3DGS-based SLAM systems while utilizing only RGB data. It operates at 30 fps or higher, a performance matched only by Photo-SLAM and GS-ICP (see Table I). While MGSO is most similar to Photo-SLAM, which combines 3DGS with ORBSLAM3, we address Photo-SLAM’s tendency to create large, memory-inefficient maps. GS-ICP offers exceptional speed but requires depth data to initialize its iterative closest point tracking, whereas our system operates using only RGB data. Unlike IG-SLAM [sarikamis2024igslam], which uses pseudo-depth RGB-D data at the cost of slower performance, MGSO maintains real-time speeds while generating accurate, compact maps.
@@ III-A SLAM module:
The tracking backbone of our system is built upon a lineage of visual SLAM approaches originating from Direct Sparse Odometry (DSO) [dso:engel18]. DSO’s key innovation is demonstrating that selective pixel sampling for photometric tracking enhances localization accuracy compared to using all pixels in an image. We chose to build upon DSO because we found that its pixel selection strategy also aligns well with initializing 3DGS. DSO tracks a set of pixels across consecutive frames and , optimizing the camera pose () by minimizing the photometric loss equation below for each pixel tracked,
@@
We observed that 3DGS performs better with slightly misplaced initialized points in flat areas than with none at all. Therefore, we implemented an interpolation method that estimates point locations in low-gradient regions based on nearby tracked pixels. Our method employs the Delaunay triangulation algorithm [LAWSON1972365] to divide the image into a series of triangles using tracked pixels as vertices. The depth of each interpolated point is calculated as the average depth of the triangle’s vertices, which generally provides accurate results for pixels on flat surfaces. While feature-based systems like Photo-SLAM also interpolate inactive 2D feature points, our method outperforms theirs due to a higher initial point count and by focusing interpolation on flat areas where it’s most accurate, which can be observed in Figure 1.
@@ III-B Dense Reconstruction
MGSO employs 3DGS as its dense reconstruction method. Following the original 3DGS, we map the scene using a set of anisotropic Gaussians . Each Gaussian is modeled with an opacity, rotation, location, scale, and color. We follow Mono-GS’ [Matsuki:Murai:etal:CVPR2024] technique of representing color using RGB instead of spherical harmonics because Mono-GS showed this increased speed for minimal impact on reconstruction quality. We render RGB images of the map using the original differentiable tile-based rasterization introduced in 3DGS.
@@
In order to improve the speed of our system, we employ Gaussian-pyramid based learning introduced in Photo-SLAM [hhuang2024photoslam] to progressively train the Gaussian map. The pyramid helps accelerate training for live video scenarios. A multi-scale Gaussian pyramid is created by repeatedly smoothing and down-sampling ground-truth image captured by the camera. The photometric loss calculation progresses from using the highest pyramid level for in initial iterations to lower levels as training advances. Furthermore, we use an optimized version of the 3DGS CUDA back-end [Patas] that is faster than the original.
@@
However, we noticed the emergence of floaters when utilizing the 3DGS’s densification and pruning strategies. To mitigate the presence of floaters, we utilized the adaptive control strategies in AbsGS [ye2024absgs]. Thus, as part of our adaptive control strategy we periodically densify Gaussians with high homo-directional view-space position gradients by splitting or cloning them. Large, high-variance Gaussians are split, while small Gaussians in under-reconstructed regions are cloned. Furthermore, we periodically prune Gaussians with low opacity to remove transparent floaters. We use the same splitting and cloning parameters as original 3DGS.
@@ IV Experiments and Discussions
We evaluate MGSO against the latest state-of-the-art 3DGS dense SLAM systems: MonoGS [Matsuki:Murai:etal:CVPR2024], GlORIE-SLAM [sandström2024splatslamgloballyoptimizedrgbonly], Splat-SLAM [keetha2024splatam], IG-SLAM [sarikamis2024igslam], and Photo-SLAM [hhuang2024photoslam], to demonstrate our system’s combination of high-quality reconstruction, efficient runtime, and compact maps.
@@ IV-A Implementation and Setup
Datasets: Evaluations are done on the sythetic Replica [replica19arxiv] dataset and real-life EuRoC MAV [doi:10.1177/0278364915620033] and TUM-RGBD [sturm12iros] datasets. These datasets are commonly used to evaluate dense SLAM systems.
@@
Experimental Setup: We utilize the default optimization configuration of 3DGS with the exception of adjusting the densification interval to 1000. We configure the SLAM module (DSO [dso:engel18]) to the default tracking settings. The parameters for increasing SLAM output point density through the inclusion of untracked points were determined through iterative testing.
@@
Evaluation: To ensure a fair comparison, we evaluated our system by inputting the output 3DGS maps and pose estimation data into the original 3DGS rendering and metric scripts. We evaluate our reconstructions with the standard image quality metrics: PSNR, SSIM [1284395], and LPIPS [zhang2018perceptual]. Using an third-party evaluation system rather than built-in metrics offers a more realistic assessment, accounting for real-world factors like potential misalignment between the poses and map. Consistent with other dense SLAM systems, we evaluate on every fifth frame. Photo-SLAM’s evaluations are updated using this methodology to ensure consistency. We did ten runs for the Replica and EuRoC dataset and five runs for the TUM-RGBD dataset. Results for other systems were obtained from their respective publications, with the exception of Mono-GS, whose results were sourced from Splat-SLAM [sandström2024splatslamgloballyoptimizedrgbonly].